OpenStudy (anonymous):

Solve Each Equation (3x-5)(2x+9)=0

OpenStudy (unklerhaukus):

a product of two term is zero, therefore one of the terms must be zero either (3x-5) =0 or (2x+9)=0

OpenStudy (unklerhaukus):

if x=5/3 the first product is zero

OpenStudy (anonymous):

3x-5=0 3x=5 x=5/3 or 1.66666 answer. 2x+9=0 2x=-9 x=-9/2 or -4.5 answer.